### v1

@@ 4.2 Artifact Design
RAG is the core framework driving the system, seamlessly combining advanced information retrieval techniques with the generative capabilities of LLMs. To support this, the system starts by scraping feedback data from the Have Your Say platform. This feedback, often unstructured and multilingual, is preprocessed to remove noise and inconsistencies. The data is then transformed into dense vector embeddings using OpenAI’s text-embedding-3-small [^5] https://platform.openai.com/docs/guides/embeddings, which encode the semantic meaning of the text into a high-dimensional numerical representation. These embeddings are stored in a MongoDB Atlas vector database, enabling efficient semantic search during retrieval.
When a user submits a query, the system first processes it by generating a vector embedding for the input, matching it against the database of stored feedback using cosine similarity. The system then returns the K entries from the database with the highest cosine similarity as the context for LLM. At this stage, the system leverages GPT-4o-mini by feeding the retrieved context along with the user’s query into the LLM. This integration delivers reliable, insightful summaries and actionable recommendations in an intuitive, human-like format. Additionally, a microphone icon allows users to input queries via voice, enhancing accessibility and ease of use. The “Ask” button initiates the query process, leading users to an analysis page that displays summaries and categorized insights derived from the data. To increase user confidence and transparency, the system includes a Source section, which highlights the resources used to generate the answer to the given query. A new chat icon positioned beside the search bar enables users to refresh the page and initiate a new query.
The output of the chatbot is structured into three parts. The first part provides an overview, highlighting key themes and sentiments from public opinions. The second part offers categorized insights, showcasing feedback segmented by different user groups and summarizing key points for each group. Finally, the third part delivers actionable insights, presenting two to three specific, evidence-based recommendations or highlighting conflicts that require resolution.
@@ 4.3 Demonstration and Evaluation
A total of 357 in-vivo codes were derived from the 11 semi-structured interviews, ranging between 44 and 74 minutes. These were then clustered into the solution objective categories, namely, real-time interaction (85 in-vivo codes), structured feedback management (69 in-vivo codes), and multilingual synthesis (63 in-vivo codes). A fourth category, prospective stakeholder usage, was introduced to inform future research directions, covering the remaining 140 in-vivo codes. The analysis shows that the participants shared a rather positive perception of the EU, with an average of 8.0 on a scale from 1 (very negative) to 10 (very positive). When asked to assess the degree of impact the EU has on their daily lives, with 1 being ’nonexistent’ to 10 being ’omnipresent’, responses averaged 7.7. In the following, the results of the qualitative assessment are presented in accordance with the identified solution objectives.
@@
Real-Time Interaction: While some participants were critical of the actual benefit over established chatbots (ID4-6), as exemplified by ID5 (“Everything sounds reasonable. I probably would need to ask the same question to ChatGPT, because, based on what I read, though it is neither right nor wrong, I do not see why we need another chatbot.”), the majority of the participants (ID 1-2,7-9) indicated they are more likely to use AskThePublic over the Have Your Say platform due to the enhanced interaction, as exemplified by ID8 (“I would be much more likely to use it than to browse through every single initiative or directive and to read the comments separately, this is obviously made much easier. And it also reduces the entry barrier, so to say, to inform myself.”). One participant narrowed the benefit of the solution down to a single statement: “Basically, the more advanced search engine for public opinions.(ID4)” Participants have particularly noted the intuitiveness of using the chatbot: “It’s easy. In the beginning, you are doing a test. If you are not sure and then you see how it works. (ID9)” It was further noted that the tool is well suited for information purposes (ID 2-3, 6-10), particularly in cases where one does not have pre-existing knowledge of a domain: “So, if you do not have a very precise question in mind, but you rather want to assess opinion in certain areas of topics, then, it is being summarized really well. […] I really like how it works because you can receive various opinions in a fairly well-structured and simple way. (ID3)” Moreover, several participants (ID 1-2,7,10-11) noted the various features of the chatbot positively: “This is actually something I really enjoy, that I have my text on the left side, and on the right side the sources for validation. (ID10)”
@@
Structured Feedback Management: Participants generally view the structure favorably (ID 1,3-4,9-11), as exemplified by ID10: “I actually like that it gives a broad answer at the top, then going into more insights and also having an implication section. So it has this tunneling effect.” The structure allows its users to conduct further research: “It is not too long, not too short, so it provides you an overview and from there you can take it further and research yourself. (ID9)” Participants particularly highlighted the ability to engage with specific stakeholder groups directly (ID1-5, 7-8, 10-11): “If you ask specific groups of people questions, it seems that you can engage with these people. You can engage with the answers that they have already provided on the platform. (ID7)” At the same time, through RAG-based computation, AskThePublic only retrieves stakeholder feedback that relates directly to the query, in stark contrast to the Have Your Say platform. While this was noted positively, all participants questioned how the sources are being selected (ID1-11): “If you say, 1000 people have responded, but now I get four random universities as output, I question the integrity of the process, what makes them stand out (ID6)?” Participants noted an apparent lack of credibility when it comes to certain stakeholders that are being displayed: “Perhaps the German AI Association I have heard before, but I do not know any other, so I cannot judge if that is a source where perhaps three people are sitting who simply do not have any relevance, or if it is some sort of NGO which is truly important. (ID3)” One participant suspected a potential bias towards more polarizing opinions that are selected to be displayed: “I think negative opinions are much more polarizing. […] So for example, not everything about dynamic tariffs is bad, but it feels like this kind of overview right now focuses pretty heavily on the bad side. (ID11)” When being asked very specific questions, the chatbot was found to create answers artificially. This is recognized as a major threat to the validity: “So if you go and search yourself, and you find out that there is nothing, this is very bad for the reputation of the system. (ID9)”
@@
Multilingual Synthesis: The solution offers the translation of stakeholder feedback given in any official EU language to the language the question was asked in, making it easier for EU citizens to understand the feedback (ID1, 10-11), as exemplified by ID 10: “The generation of my grandparents, or even my parents, they do not speak English so well. So for them, this is valuable because if they are struggling to understand the text, they will go to some tabloid newspaper, even if they do not provide accurate information.” However, it was noted that this only applies to the immediate text output from the chatbot and does not apply to the sources panel. Further, finding the right sourcing balance is commonly criticized (ID4-5, 10-11). On the one hand, participants found larger countries with bigger impact most relevant, as exemplified by ID 4 (“In this question, I only got answers from Germany and Slovakia. Germany, I understand, but as for Slovakia, frankly, I do not understand why it is being displayed.”), on the other hand, participants noted that they would like to have a bigger variety, as exemplified by ID 10 (“I would prefer to have three bullet points from three different countries instead of having the majority from one country.”).
@@ 5 Discussion
Utilizing DSR, we are able to answer our RQ: How can e-participation data be leveraged by LLMs to support trust in policy-making processes within the EU? We do so by discussing the solution objectives and design features of the designed artifact AskThePublic (Section 5.1). Further, we give insights on how different stakeholders can utilize the given design (Section 5.2).
5.1 Designed Artifact
@@
The evaluation of the designed artifact demonstrated that each of our three solution objectives holds significant promise for enhancing e-participation.
@@
Real-time interaction: Participants appreciated the chatbot’s intuitiveness and ease of use, noting that its real-time interaction capabilities significantly enhanced user engagement compared to traditional platforms. This finding aligns with the MRT, which posits that richer media facilitate more effective communication [11, 12]. The chatbot’s ability to provide immediate, interactive responses mirrors the high richness of face-to-face communication, thereby increasing its effectiveness in engaging users and being a practical example of conversational government [7, 4].
@@
Structured Feedback Management: AskThePublic’s ability to organize unstructured feedback into coherent and actionable insights was another key strength noted by participants. By leveraging RAG, the system effectively filters and structures feedback, ensuring that only relevant information is presented [26], thus focusing on transparency, a quality needed in e-participation systems [3, 23]. While this concept was rooted in Sprenkamp et al., [41], who analyzed social media data to make it available for policy-making, AskThePublic allows for generating insights from e-participation data in natural language, compared to prior ML-based tools [6, 21, 29]. As participants found this structured format facilitated easier navigation and deeper understanding of the e-participation data, we foresee that systems like AskThePublic will attract higher adaptation rates, thus addressing a common problem among modern e-participation tools [45, 21]. Although most of the feedback was positive, a participant questioned whether a GPT model without RAG could produce similar responses. However, without the public context provided by RAG, its performance would suffer, provided there is relevant data to display. Thus, this comment underlines that the system is only as good as the available data.
@@
Multilingual Synthesis: The multilingual capabilities of AskThePublic were instrumental in making feedback accessible across the EU’s 24 official languages. This feature directly addresses language barriers identified as significant determinants affecting e-participation [3]. By utilizing LLMs for translation and synthesis, the chatbot fosters a more inclusive environment, aligning with the principles of conversational government that advocate for language-inclusive communication channels [7]. Further, while the chatbot answers in any of the official EU languages, it further synthesizes context from various source documents in the RAG-system [41]. However, participants noted that while the chatbot translated immediate text outputs, the sources panel remained inaccessible in non-English languages. This limitation highlights the need for comprehensive multilingual support to fully realize inclusivity [31]. Additionally, concerns regarding the balance and diversity of sources were raised, indicating that the chatbot should implement strategies to ensure a more equitable distribution of feedback from various countries. Notably, the current embedding fetching mechanism prioritizes data points in the language of the query, which may inadvertently limit the diversity of retrieved multilingual feedback, highlighting the need to balance language prioritization within the retrieval process.
@@
In addition, users proposed new features for AskThePublic. To mitigate the bias towards polarizing opinions and ensure a more balanced representation of feedback, the system could incorporate algorithms that promote diversity in source selection, aligning with research on unbiased data representation in public administration [3]. Another option is to utilize custom models based on open-source projects. Currently, we utilize the GPT API, a closed-source model. Thus, the training data and potential biases cannot be identified; here, open-source LLMs like EURO-LLM [30] could be leveraged to align more closely with EU values of transparency and data sovereignty. While open-source models often lack the accuracy of closed-source models, this has been recently challenged by the development of DeepSeek [18], which exceeds the efficiency of state-of-the-art LLMs while being on par in accuracy. Thus, we expect governments to be empowered in the future to utilize open-source solutions. Several users flagged that AskThePublic could integrate unique functionalities not offered by existing chatbots like ChatGPT, such as specialized data visualization tools or tailored stakeholder engagement features, enhancing its distinct value proposition [8]. Furthermore, expanding the chatbot’s integration capabilities with other e-participation platforms and social media channels can broaden its accessibility and usability, as highlighted in related works on digital engagement strategies [28]. By implementing these recommendations, future iterations of AskThePublic can enhance its effectiveness, inclusivity, and trust, thereby further supporting democratic engagement within the EU.
@@ 5.2 Possible User Groups
Tool for EU Policymakers: EU policymakers can leverage AskThePublic to gather and analyze feedback more effectively, thus informing more inclusive policy decisions, which are based on public data. A participant expressed concern about the representativeness of feedback: “I would not like to create a statistic out of the respondents from the platform right now, because I think it is not representative. (ID7)” By incorporating features that balance data distribution across different demographics, the chatbot can enhance both the quality and reliability of the feedback used in policy-making, addressing challenges related to sampling bias and data integrity [3, 23].
Tool for EU Citizen: The AskThePublic chatbot holds significant potential for politically engaged EU citizens by providing an intuitive and interactive platform for accessing and analyzing policy feedback. One participant remarked: “It serves its purpose if I am interested to learn, then it is a very good tool. It helps me at navigating all the different websites. (ID2)” Other participants valued the empowerment and active participation the tool could foster (ID8, ID6, ID10). By making feedback more accessible and structured, the chatbot can enhance citizen engagement and address common barriers to e-participation, such as lack of information and perceived impact.
Tool for Journalists and Researchers: For journalists and researchers, AskThePublic provides a valuable tool for extracting feedback on legislation. Thus, they can identify the sentiments of citizens or organizations towards a given topic. One participant underscored the importance of source verification: “In the main text I cannot see a match with the source. It would be very helpful to see that this is taken from source number 7, for example. (ID10)” Ensuring clear links to original feedback can support accurate reporting and data-driven research, aligning with best practices in transparency and reliability.
@@ 6 Conclusion
While e-participation has been in the focus of the academic discourse [6, 21, 29], methods to analyze the submitted feedback are sparse [20, 2, 5], focusing on the fragmented usage of ML algorithms for the classification of citizen needs. The advent of LLMs gives the possibility to create applications that utilize natural language for analyzing citizen feedback, needing little to no understanding of the technology as shown through AskThePublic.
Utilizing the DSR [22, 33], the development of AskThePublic has made three contributions. First, we give a first tool for the active analysis of e-participation using LLMs, enabling users to ask questions instead of needing expertise in ML or any computational tool, which is open to the general public. Second, we give design recommendations for tools enabling the structured analysis of e-participation. We hope that these recommendations are useful for further development in academia and the public sector. Third, we identify first user groups for applications like AskThePublic. However, our study is not without limitations. Our evaluation was solely done with a rather small user group with largely positive views of the EU, leaving room for further iterations of DSR and larger evaluations. Moreover, we developed and tested AskThePublic solely using the GPT-4 API, creating a bias towards the training data of this closed-source model.
We see these limitations as opportunities for future work. Later, iterations of AskThePublic or similar tools should further evaluate our design recommendations with the goal of generating long-lasting design principles [17]. Moreover, the effect of different closed and open-source LLMs should be tested to align with the values of the EU, e.g., EURO-LLM [30].
{credits}

### v2

@@ 4.2 Artifact Design
RAG is the core framework driving the system, combining advanced information retrieval techniques with the generative capabilities of LLMs. We rely on RAG because even models like GPT-4 can produce inaccurate information, and RAG has proven effective in mitigating this issue. To support this, the system starts by scraping feedback data from the Have Your Say platform. This feedback, often unstructured and multilingual, is preprocessed to remove noise and inconsistencies. The data is then transformed into dense vector embeddings using OpenAI’s text-embedding-3-small, which encode the semantic meaning of the text into a high-dimensional numerical representation. These embeddings are stored in a MongoDB Atlas vector database, enabling efficient semantic search during retrieval.
When a user submits a query, the system first processes it by generating a vector embedding for the input, matching it against the database of stored feedback using cosine similarity. The system then returns the $K$ entries from the database with the highest cosine similarity as the context for LLM. At this stage, the system leverages GPT-4o-mini by feeding the retrieved context along with the user’s query into the LLM. This integration delivers reliable, insightful summaries and actionable recommendations in an intuitive, human-like format. The RAG pipeline process to generate final response is shown in Figure 2.
The output of the chatbot is structured into three parts. The first part provides an overview, highlighting key themes and sentiments from public opinions. The second part offers categorized insights, showcasing feedback segmented by different user groups and summarizing key points for each group. Finally, the third part delivers actionable insights, presenting two to three specific, evidence-based recommendations or highlighting conflicts that require resolution. To increase user confidence and transparency, the system includes a Source section, which highlights the resources used to generate the answer to the given query. A new chat icon positioned beside the search bar enables users to refresh the page and initiate a new query.
@@ 4.3 Demonstration and Evaluation
A total of 357 in-vivo codes were derived during the analysis of the 11 semi-structured interviews. These were then clustered into the solution objective categories, namely, real-time interaction (85 in-vivo codes), structured feedback management (69 in-vivo codes), and multilingual synthesis (63 in-vivo codes). A fourth category, prospective stakeholder usage, was introduced to inform future research directions, covering the remaining 140 in-vivo codes. The analysis of the pre-study shows that the participants shared a rather positive perception of the EU, with an average of 8.0 on a scale from 1 (very negative) to 10 (very positive). In the following, the results of the qualitative assessment according to the solution objectives are presented.
@@
Real-Time Interaction Although some participants (ID4-6) were sceptical whether the RAG capabilities provide enough benefit over established chatbots (“Everything sounds reasonable. I probably would need to ask the same question to ChatGPT, because, based on what I read, though it is neither right nor wrong, I do not see why we need another chatbot.”, ID5), the majority of the participants (ID 1-2,7-9) indicated they are more likely to use AskThePublic over the Have Your Say platform due to the enhanced interaction: “I would be much more likely to use it than to browse through every single initiative or directive and to read the comments separately, this is obviously made much easier. And it also reduces the entry barrier, so to say, to inform myself. (ID8)” One participant narrowed the benefit of the solution down to a single statement: “Basically, the more advanced search engine for public opinions. (ID4)” Participants have particularly praised its intuitiveness: “It’s easy. In the beginning, you are doing a test. If you are not sure, then you see how it works. (ID9)” The tool was further noted to be well suited for information purposes (ID 2-3, 6-10), particularly in cases where one does not have pre-existing knowledge of a domain: “So, if you do not have a very precise question in mind, but you rather want to assess opinion in certain areas of topics, then, it is being summarized really well. […] I really like how it works because you can receive various opinions in a fairly well-structured and simple way. (ID3)” Moreover, several participants (ID 1-2,7,10-11) noted the various features of the chatbot positively: “This is something I really enjoy, that I have my text on the left side, and on the right side the sources for validation. (ID10)
@@
Structured Feedback Management Participants generally view the structure favorably (ID 1,3-4,9-11), as exemplified by ID10: “I actually like that it gives a broad answer at the top, then going into more insights and also having an implication section. So it has this tunneling effect.” The structure allows its users to conduct further research: “It is not too long, not too short, so it provides you an overview and from there you can take it further and research yourself. (ID9)” Participants particularly highlighted the ability to engage with specific stakeholder groups directly (ID1-5, 7-8, 10-11): “If you ask specific groups of people questions, it seems that you can engage with these people. You can engage with the answers that they have already provided on the platform. (ID7)” At the same time, through RAG-based computation, AskThePublic only retrieves stakeholder feedback that relates directly to the query, in stark contrast to the Have Your Say platform. While this was noted positively, all participants questioned how the sources are being selected (ID1-11): “If you say, 1000 people have responded, but now I get four random universities as output, I question the integrity of the process, what makes them stand out (ID6)?” Participants noted an apparent lack of credibility when it comes to certain stakeholders that are being displayed: “Perhaps the German AI Association I have heard before, but I do not know any other, so I cannot judge if that is a source where perhaps three people are sitting who simply do not have any relevance, or if it is some sort of NGO which is truly important. (ID3)” One participant suspected a potential bias towards more polarizing opinions that are selected to be displayed: “So for example, not everything about dynamic tariffs is bad, but it feels like this kind of overview right now focuses pretty heavily on the bad side. (ID11)” When being asked very specific questions, the chatbot was found to create answers artificially. This is recognized as a major threat to the validity: “So if you go and search yourself, and you find out that there is nothing, this is very bad for the reputation of the system. (ID9)”
@@
Multilingual Synthesis The solution translates stakeholder feedback given in any official EU language to the language the question was asked in, making it easier for EU citizens to understand the feedback (ID1, 10-11), as exemplified by ID 10: “The generation of my grandparents, or even my parents, they do not speak English so well. So for them, this is valuable because if they are struggling to understand the text, they will go to some tabloid newspaper, even if they do not provide accurate information.” However, it was noted that this only applies to the immediate text output from the chatbot and does not apply to the sources panel. Further, finding the right sourcing balance is commonly criticized (ID4-5, 10-11). On the one hand, participants found larger countries with bigger impact most relevant, as exemplified by ID 4 (“In this question, I only got answers from Germany and Slovakia. Germany, I understand, but as for Slovakia, frankly, I do not understand why it is being displayed”), on the other hand, participants noted that they would like to have a bigger variety, as exemplified by ID 10 (“I would prefer to have three bullet points from three different countries instead of having the majority from one country.”)
@@ 5 Discussion
Utilizing DSR, we are able to answer our RQ. We do so by discussing the solution objectives and design features of the designed artifact AskThePublic (Section 5.1). Further, we give insights on how different stakeholders can utilize the given design (Section 5.2).
5.1 Designed Artifact and Future Research
@@
The evaluation of the designed artifact demonstrates that each of our three solution objectives holds significant promise for enhancing e-participation.
@@
Participants appreciated the chatbot’s intuitiveness and ease of use, noting that its real-time interaction capabilities significantly enhanced user engagement compared to traditional platforms. This finding aligns with the MRT, which posits that richer media facilitate more effective communication ([56]. The chatbot’s ability to provide immediate, interactive responses mirrors the high richness of face-to-face communication, thereby increasing its effectiveness in engaging users as a practical example of conversational government [51].
@@
AskThePublic’s ability to organize unstructured feedback into coherent and actionable insights was another key strength noted by participants. However, in the absence of relevant information to display, the chatbot might hallucinate, thus leading some participants to question the added benefits over LLMs without RAG, underlining that the system is only as good as the available data. Provided there is relevant data, the system effectively filters and structures feedback, ensuring that only relevant information is presented [68], thus focusing on transparency, a quality needed in e-participation systems [47]. While this concept was rooted in [*]sprenkamp2025data (2025), who analyzed social media data to make it available for policy-making, AskThePublic allows for generating insights from e-participation data in natural language, a large contribution over prior ML-based tools [50, 64, 71].As participants found this structured format facilitated easier navigation and deeper understanding of the e-participation data, we foresee that systems like AskThePublic will attract higher adaptation rates, thus addressing a common problem among modern e-participation tools [87, 64].
@@
The multilingual capabilities of AskThePublic helped make feedback accessible across the EU’s 24 official languages. This feature directly addresses language barriers identified as significant determinants affecting e-participation [47]. By utilizing LLMs for translation and synthesis, the chatbot fosters a more inclusive environment, aligning with the principles of conversational government that advocate for language-inclusive communication channels [51]. Further, while the chatbot answers in any of the official EU languages, it further synthesizes context from various source documents in the RAG-system [84]. Concerns regarding the balance and diversity of sources were raised, indicating that the chatbot should implement strategies to ensure a more equitable distribution of feedback from various countries. Notably, the current embedding fetching mechanism prioritizes data points in the language of the query, which may inadvertently limit the diversity of retrieved multilingual feedback, highlighting the need to balance language prioritization within the retrieval process. These limitations highlight the need for comprehensive multilingual support to further inclusion.
@@
In addition, users proposed new features for AskThePublic. To mitigate the bias towards polarizing opinions and ensure a more balanced representation of feedback, the system could incorporate algorithms that promote diversity in source selection, aligning with research on unbiased data representation in public administration [47]. Another option is to utilize custom models based on open-source projects. Currently, we utilize the GPT API, a closed-source model. Thus, the training data and potential biases cannot be identified; here, open-source LLMs like EURO-LLM [72] could be leveraged to align more closely with EU values of transparency and data sovereignty. While open-source models often lack the accuracy of closed-source models, this has been recently challenged by the development of DeepSeek [61], which exceeds the efficiency of state-of-the-art LLMs while being on par in accuracy. Thus, we expect governments to be empowered in the future to utilize open-source solutions. Several users flagged that AskThePublic could integrate unique functionalities not offered by existing chatbots like ChatGPT, such as specialized data visualization tools or tailored stakeholder engagement features, enhancing its distinct value proposition [53]. Furthermore, expanding the chatbot’s integration capabilities with other e-participation platforms and social media channels can broaden its accessibility and usability, as highlighted in related works on digital engagement strategies [70]. By implementing these recommendations, future iterations of AskThePublic can enhance its effectiveness, inclusivity, and trust, thereby further supporting democratic engagement.
Last, in line with MRT [56, 57], communication media that allow immediate feedback and multiple cues are more effective at reducing equivocality. AskThePublic’s chat interface, by offering real-time, two-way dialogue and transparent sourcing, functions as a richer medium compared to the static Have Your Say portal. This increased richness directly maps to participants’ reports of greater engagement and clearer understanding.
@@ 5.2 Possible User Groups
EU policymakers can leverage AskThePublic to gather and analyze feedback more effectively, thus informing more inclusive policy decisions, which are based on public data. By incorporating features that balance data distribution across different demographics, the chatbot can enhance both the quality and reliability of the feedback used in policy-making, addressing challenges related to sampling bias and data integrity [47].
For politically engaged EU citizens, AskThePublic holds significant potential by providing an intuitive and interactive platform for accessing and analyzing policy feedback. Participants further valued the empowerment and active participation the tool could foster (ID8, ID6, ID10). By making feedback more accessible and structured, the chatbot can enhance citizen engagement and address common barriers to e-participation, such as lack of information and perceived impact.
For journalists and researchers, AskThePublic provides a valuable tool for extracting feedback on legislation. Thus, they can identify the sentiments of citizens or organizations towards a given topic, providing researchers and journalists with a reliable and unique source that can be used for publications. However, future iterations of AskThePublic need to design clear links to original feedback that can support accurate reporting and data-driven research.
@@ 6 Conclusion
While e-participation has been in the focus of the academic discourse [50, 64, 71], methods to analyze the submitted feedback are sparse and focusing on the fragmented usage of ML methods to classify citizen needs [63, 46, 49]. The advent of LLMs enables novel applications that utilize NLP to analyze citizen feedback, needing little to no understanding of the technology, as shown through AskThePublic.
Using the DSR methodology [65, 76], the development of AskThePublic has made three contributions. First, we provide an initial tool for the analysis of citizen feedback using LLMs, which are open to the general public, enabling users to ask questions without the need for expertise in ML or computational tools. Second, we give design recommendations for tools enabling the structured analysis of e-participation. We hope that these recommendations are useful for further development in academia and the public sector. Third, we identify the first user groups for our tool.
However, our study is not without limitations. Our evaluation was solely done with a rather small user group with largely positive views of the EU, leaving room for broader evaluations. Moreover, we developed and tested AskThePublic solely using the GPT-4 API, creating a bias towards the training data of this closed-source model. We see these limitations as opportunities for future work. Later, iterations of AskThePublic or similar tools should further evaluate our design recommendations with the goal of generating long-lasting design principles [60]. Moreover, the effect of different closed and open-source LLMs should be tested to align with the values of the EU, e.g., EURO-LLM [72].
Acknowledgements This research was funded in part by the Luxembourg National Research Fund (FNR) and PayPal, PEARL grant reference 13342933/Gilbert Fridgen. We also thank the University of Zurich and the Digital Society Initiative for (partially) financing the study through Liudmila Zavolokina’s DIZH postdoc fellowship. For the purpose of open access, and in fulfillment of the obligations arising from the grant agreement, the author has applied a Creative Commons Attribution 4.0 International (CC BY 4.0) license to any Author Accepted Manuscript version arising from this submission.